OpenStudy (anonymous):

Stephen earned scores of 88 points, 94 points, 89 points, and 85 points on four tests. What is the lowest score he can get on the fifth test and still finish with an average score of 90 points? 94 points 95 points 96 points 97 points

OpenStudy (anonymous):

my answer is 96.. is that right?

OpenStudy (anonymous):

\[\frac{ 88 + 94 + 89 + 85 + x }{ 5 } = 90\] \[\frac{ 356 + x }{ 5 } = 90\]

OpenStudy (anonymous):

so.. is that a yes? or..

OpenStudy (anonymous):

356 + x = 450 x = 94

OpenStudy (anonymous):

Yes, it is A.